Problem 2: Using a Spreadsheet to Calculate Mortgage Payments
What is the monthly payment on a $150,000, 30-year mortgage if the mortgage rate is 5.75 percent? 6.25 percent?
7.5 percent? 9 percent?
Present Annual Periodic => Payment
value Periods Interest Rate Interest Rate
$150,000 FORMULA 5.75% FORMULA = FORMULA
$150,000 6.25% =
$150,000 7.50% =
$150,000 9.00% =
Answer #1

Monthly installment payment amount for mortgage is given by (M):

M = (P * r) / [1 - (1+r)-n ]

where P = mortgage amount, r= monthly interest rate = annual interest rate / 12, n = number of monthly installments = 30*12 = 360

A B C D E F 13 => Payment 16 14 Present 15 value $150,000 17 $150,000 18 $150,000 19 $150,000 Periods 360 360 360 360 | Annua

The excel cell with formula for the above sheet is copied below:

Present Annual Periodic => Payment
value Periods Interest Rate Interest Rate
150000 =30*12 0.0575 =C16/12 = =D16*A16/(1-(1+D16)^-B16)
150000 =30*12 0.0625 =C17/12 = =D17*A17/(1-(1+D17)^-B17)
150000 =30*12 0.075 =C18/12 = =D18*A18/(1-(1+D18)^-B18)
150000 =30*12 0.09 =C19/12 = =D19*A19/(1-(1+D19)^-B19)